OpenStudy (anonymous):

Solve the inequality 3(-x+2)<27 A. X>7 B. x>-7 C. X<7 D. X<-7

OpenStudy (anonymous):

@angela210793

OpenStudy (anonymous):

@Jfm

OpenStudy (anonymous):

its d) x < -7

OpenStudy (anonymous):

@Jfm identify the slope in this linear equation 3x+y=6 A. 6 B. 2 C. -3 D. 3

OpenStudy (anonymous):

its c) -3

OpenStudy (anonymous):

How is it negative? @Jfm

OpenStudy (anonymous):

because you subtract 3x to the other side and since the linear equation is y=mx+b, whatever is in the m position is youre slope

OpenStudy (anonymous):

@Jfm if n=2 and x=1/2 then n(4-x)= A. 7 B C. 10 D. 1

OpenStudy (anonymous):

B. 3

OpenStudy (anonymous):

the answer is a) 7

jhonyy9 (jhonyy9):

so there is -3x+6<27 -3x<21 divide by -3 and will get x>21/(-3) x>-7

OpenStudy (anonymous):

@Jfm solve the system of equations Y=x-4 Y=-x+2 A. (1,2). B. (3,-1) C. (0,3) D. (-1,3)

OpenStudy (anonymous):

its b) (3,-1)

OpenStudy (anonymous):

@jhonyy9 identify the y-intercept in this linear equation 3x+y=6 A. 2 B. 3. C. -3. D. 6

OpenStudy (anonymous):

its d) 6

jhonyy9 (jhonyy9):

jfm your answer on the first exercise not is right please you check it the right answer is B

jhonyy9 (jhonyy9):

because when you divide 21 by -3 so this is equal -7 and not 7

OpenStudy (anonymous):

right i did say it was equal to -7. i think you mean i forgot to flip the sign

OpenStudy (anonymous):

@jhonyy9 identify the y-intercept in this linear equation 3x+y=6 A. 2 b. 3 c. -3 d. 6

jhonyy9 (jhonyy9):

ok so what is your opinion ?

jhonyy9 (jhonyy9):

how you begin it ?

jhonyy9 (jhonyy9):

add to both sides -3x ok ?

OpenStudy (anonymous):

right